# Forecasting the future development in quality and value of professional football players

### v2

@@ Abstract
Transfers in professional football (soccer) are risky investments because of the large transfer fees and high risks involved. Although data-driven models can be used to improve transfer decisions, existing models focus on describing players’ historical progress, leaving their future performance unknown. Moreover, recent developments have called for the use of explainable models combined with uncertainty quantification of predictions. This paper assesses explainable machine learning models based on predictive accuracy and uncertainty quantification methods for the prediction of the future development in quality and transfer value of professional football players. The predictive accuracy is studied by training the models to predict the quality and value of players one year ahead. This is carried out by training them on two data sets containing data-driven indicators describing the player quality and player value in historical settings. In general, the random forest model is found to be the most suitable model because it provides accurate predictions as well as an uncertainty quantification method that naturally arises from the bagging procedure of the random forest model. Additionally, this research shows that the development of player performance contains nonlinear patterns and interactions between variables, and that time series information can provide useful information for the modeling of player performance metrics. The resulting models can help football clubs make more informed, data-driven transfer decisions by forecasting player quality and transfer value.
@@
Football analytics, football scouting, explainable machine learning, player quality, player value, practitioner-oriented research, player development prediction, uncertainty quantification
@@ 1 Introduction
Transfers in professional football (soccer) are a risky business because the average transfer fee has increased in recent years [1] and because these fees can be characterized as investments with high risks where large fees are involved [2]. Extensive knowledge about players is beneficial in making well-informed decisions about these complex transfer investments in football. This paper shows how football players’ development in quality and monetary value can be predicted to enhance decision-making.
@@
Models providing information about player quality and value have recently emerged with the evolution of data-driven player performance indicators. Improvements in data-capturing technologies resulted in large data sets containing in-game data about football players, which provide the opportunity to obtain more complex variables on player performance [3, 4]. Numerous player performance indicators have been introduced since then. An example is the expected goals (xG) indicator, which values shot chances and shooting ability [5, 6, 7, 8]. Next to such action-specific models, assessment methods exist for general player performance, which can be divided into bottom-up and top-down ratings [9]. Expected threat (xThreat) [10, 11, 12] and VAEP [13, 14, 15, 16] are examples of bottom-up ratings that quantify action quality and use the quality of the actions to create general ratings. Top-down ratings such as plus-minus ratings [17, 18, 19, 20, 21], Elo ratings adjusted for team sports [22], and SciSkill algorithm [23] distribute credit of player performance based on the result of a team as a whole. For the monetary value of players, many models about the estimation of transfer fees and market values have been introduced and provide indicators of the current value of football players [24]. These performance and financial models describe the quality and monetary value of a football player, and these can complement traditional scouting reports. This allows managers and technical directors of football clubs to make better-informed transfer decisions.
@@
These models for the player quality and the transfer value give information about the quality and financial value of football players up to that moment, although a transfer decision regards whether a football player should be part of a team in the future. To make better-informed transfer decisions, team managers and technical directors also need insights into the development of the indicator values that describe the financial value and player performance in the future. This paper examines the training of supervised learning models that forecast the development in player quality and transfer value of players one year ahead.
@@
To this end, two prediction problems are studied: forecasting the quality of a football player and forecasting their transfer value one year ahead. In the first prediction problem, models are trained to predict the development of a top-down quality indicator, the SciSkill [23]. The second prediction problem concerns the prediction of the development of the player value, described by the Estimated Transfer Value (ETV) [25]. A forecasting horizon of one year is selected in this study as it aligns with the length of one full season in football. The resulting models of these prediction problems offer insight into the question of whether a player will be worth the money in the future. These models thus provide critical insights for the managerial staff of a professional football club.
@@
To further improve the usability of the research results for staff at football organizations, the findings of models should be presented such that they can be utilized by football clubs in practice, as stressed by Herold et al. [4]. This means that models should not only be assessed on predictive accuracy but also on explainability and methods for uncertainty quantification [26]. To this end, only explainable supervised learning models are used in this research, and the models are assessed on their methods for uncertainty quantification. To focus even more on the applicability for practitioners, the accuracy of the models is partly determined by estimating the loss values of different groups of football players, because certain types of football players are more important from a practitioner’s perspective.
@@
The aim of this research is to find the most suitable explainable machine learning model to forecast player performance with respect to predictive accuracy and methods for uncertainty quantification. To find this, several explainable supervised learning models are studied. Their methods for uncertainty quantification are studied based on the literature. The predictive performance is assessed by implementing them to forecast the player quality (SciSkill) and the player value (Estimated Transfer Value) of football players one year ahead. The predictive quality is studied on both the general population of players and on subgroups of players that are more interesting for practitioners, such as young or high-value players. An overview of this process is visualized in Figure 1. These results will then be combined to find an answer to the research question.
@@
The main contribution of this paper is the illustration of how extra value can be added to football player KPIs by forecasting future values. The results of this study provide knowledge on what models are most suitable for application by practitioners, lowering the threshold for real-life implementation. By taking explainability, uncertainty quantification, and performance on important subpopulations of players into account, this research is presented such that the results can be used easily by practitioners. In this way, this paper also contributes to bridging the gap between academic research and practitioners.
This paper is organized as follows: The scientific background on existing literature is given in Section 2. The long-term forecasting of player development is then studied for the two prediction problems, for which the methods are described in Section 3. The results of the models in the prediction problems of the player quality and player value are presented in Section 4. The conclusions of this research are then summarized in Section 5, followed by a discussion of the implications and future directions in Section 6.
@@ 2.1.1 Considered models
The first linear model considered is ordinary least squares (OLS), sometimes called (multiple) linear regression as described by Hastie et al. [27, Section 3.2]. This model assumes that the true relation is linear $y_{i}=X_{i}\beta+\varepsilon_{i}$, where $\varepsilon_{i}$ is normally distributed noise. OLS minimizes the residual sum of squares
@@
and has the lowest mean square error of all unbiased estimators. Lasso regression [28] is another linear model that introduces a bias in the model by minimizing the penalized residual sum of squares $RRS(\beta)+\lambda\sum_{i=1}^{p}|\beta_{i}|$. By introducing this bias, it is possible to provide more accurate predictions. The third linear model is the linear mixed effects (LME) model, which assumes that
@@
where $\varepsilon_{i}\sim N(0,\sigma^{2}\Lambda_{i})$[29]. The random variable $Z_{i}$ is called the random effect of group $i$ and can be used to model the influence of a specific random attribute that should not influence the predictions, e.g., nationality. Because of their linearity, these models are explainable and suitable for applications.
The CART decision tree [27, Section 9.2] is a model that recursively splits the feature space by taking splits of the form $\{X|X_{j}\leq s\},\{X|X_{j}>s\}$ such that it minimizes the sum of squares within each split. As described by Hastie et al. [27, p.312], decision trees generally suffer from a high variance. A random forest model solves this by repeatedly training decision trees on resampled data [27, Chapter 15]. The random forest model then predicts by taking the average prediction of these decision trees. Boosting is another way of improving decision trees. A boosting algorithm fits decision trees sequentially and reweighs the data points for which the model has a bad performance [27, Chapter 10]. The XGBoost [30] combines several techniques, such as regularization and feature quantile estimates, to provide such a boosting algorithm. Generally, the splitting behavior in tree-based algorithms mimics the if-then reasoning of humans. Additionally, they can provide feature importances by considering how much a specific feature improves the predictions on the training set. This makes the tree-based models explainable and thus suitable models.
The last type of model considered in this paper is the k-nearest neighbor (kNN) model. Given the features, these models search for the most similar data points in the training data set and predict the associated $y_{i}$-value by taking the average of the $k$-nearest neighbors [27, Subsection 2.3.2]. The closest neighbors can be weighted more heavily as illustrated by Dudani [31]. The ReliefF algorithm by [32] calculates feature importances by looking at the probability of having different values considering the neighbors of a data point. A downside of the kNN models is that they suffer from the curse of dimensionality as illustrated in [27, Section 2.5]. Nonetheless, the kNN models can provide examples of training data points to explain predictions. In the context of the current study, this would be players who have been in similar situations in history, which gives a very intuitive interpretation to practitioners.
@@ 2.1.2 Uncertainty quantification
Linear regression has an underlying theory that gives prediction intervals as described by Neter et al. [33]. These prediction intervals are based on assumptions that are unlikely to hold for the prediction problems in this paper, such as the normality of errors. Nonetheless, they do give an indication of the uncertainty of the prediction. Similarly, the kNN models provide the neighbors, which are a group of similar data points. Uncertainty quantification can be obtained by taking the minimal and maximal values of the dependent values within these neighbors if the number of neighbors $k$ is of significant size. Lastly, the bagging procedure of the random forest model can be utilized to obtain uncertainty quantification for the predictions, as described by Wager et al. [34]. This method utilizes the different decision trees in the random forest to quantify the uncertainty in the prediction. From these properties, it is concluded that the linear regression, random forest, and kNN-based models are favorable with respect to uncertainty quantification.
@@ 2.2 Existing indicators
In recent years, the increasing amount of available data in football has driven the introduction of methods to rate individual football players [35]. The performance of a professional football player has traditionally been determined via expert judgment based on video data and statistics describing the frequency of in-game events. Data-driven models have offered the possibility to reduce the bias in assessments and improve the consistency of the judgment of both player quality and player value, as discussed below. In this way, these models have provided new and consistent insights into the quality and monetary value of football players, which has aided in transfer decisions.
@@ 2.2.1 Player performance
The creation and the comparison of models for player performance have revealed new challenges. Because teams can have different aims in football and can apply various tactics, there does not exist a ground truth for player performance [26]. A player can, for instance, be instructed to keep the ball in possession, which leads to the player performing fewer actions that might result in scoring a goal. Because of this, there is a lack of ground truth, and various models exist that describe different aspects of the game.
@@
Some models define the players’ quality by their actions. Although action-specific models exist [5, 6, 36, 7, 8], methods to assess the quality of all types of actions have been introduced that generalize the action-specific models. These general models often define a ‘good’ action as one that increases the probability of scoring and decreases the probability of conceding a goal. The VAEP model by Decroos et al. [13] calculates the probability of scoring given the last three actions, including in-game context. Because the involved machine learning techniques are considered a black-box model by practitioners, the authors in [14, 15] introduced methods to make the VAEP model more accessible for practitioners. A comparable, more interpretable framework is the xThreat model by Rudd [10], which only considers the current situation, defined by the location of the ball-possessing player, to estimate the probabilities of the ball transitioning to somewhere else on the pitch. This is modeled by a Markov chain, which can be used to describe the probabilities of scoring before and after each action to find the quality of an action. Van Roy et al. [11] have compared the xThreat and VAEP models and have found that, although the xThreat model is more interpretable to practitioners, it can only take into account the position of an action and excludes contextual information such as the position of defenders from its model. This means that there is a trade-off between explainability and the inclusion of in-game context when choosing either VAEP or xThreat models. Van Arem and Bruinsma [12] have extended the xThreat model by including variables describing the defensive situation and height of the ball. This Extended xThreat model can take into account in-game context like the VAEP model, while maintaining explainability as an xThreat model.
@@
The bottom-up ratings like xThreat and VAEP describe the quality of a professional football player by assessing in-game on-the-ball actions. These on-the-ball actions are often offensive actions, and they are better at capturing the quality of attackers and attacking midfielders, which indicates that a bias occurs when these models are studied. This bias makes it harder to assess the quality of defensive players, and it can be reduced by using top-down models that describe player quality using the lineups and outcomes. Plus-minus ratings are such ratings that were first used in ice hockey and basketball [18], and were later applied to football by Sæbø and Hvattum [17]. For plus-minus ratings, the game is partitioned into game segments that contain the same lineups, which correspond to the data points in the data set. In this data set, the result of the game segment, the goal difference, for example, is the dependent variable. Indicators describing whether a player was active in the segment are the independent variables. Linear regression is then applied to estimate the influence of players on the results. The coefficients of the regression describe the average impact of a player on the game result and give an indicator for player performance over the period of time covered by the data. Because substitutions are infrequent in football, a game does not have many game segments with different lineups. Moreover, football is a low-scoring sport. This creates a situation where a low number of segments with limited distinction in outcomes must be used to infer player quality. To deal with this, other quantities have been used as the dependent variable to describe the result of a segment, like the expected number of goals (xG), the expected number of points (xP), and the created VAEP values by Kharrat et al. [18] and Hvattum and Gelade [9]. Pantuso and Hvattum [19] additionally showed the potential of taking age, cards, and home advantage into account, and Hvattum [20] illustrated how separate defensive and offensive ratings can be obtained. Because putting a player in the lineup is an action that a coach performs, De Bacco et al. [21] adapted the plus-minus ratings using a causal model to better describe the influence of the selection of a player. These studies show how plus-minus ratings have been adapted to the application of rating players in football.
@@
Whereas the plus-minus ratings describe the quality of a player using multiple historical games, there also exist models that determine the quality of a player after each game using the lineups and final score. Elo ratings are such ratings and were originally developed to evaluate performance in one-on-one sports. The concept was subsequently adapted to the game of football by Wolf et al. [22]. This adapted algorithm provides ratings for each individual football player and calculates the team rating via the average of players in a game weighted by the number of minutes played. The ratings are then used to predict the match outcome using a fixed logistic function, and after each game, the individual ratings are adjusted. If the outcome is better than predicted, the player rating is increased, and if the outcome is worse than expected, it is decreased. The authors also introduced an indicator for player impact to deal with the fact that this rating undervalues good players at below-average teams. The SciSkill [23] is a model that generalizes the Elo rating in several aspects. Instead of only considering one player’s quality, it describes football players using a defensive rating and an offensive rating. These offensive and defensive scores are then combined to obtain one value, the SciSkill. For each game, the outcome of the game is predicted using a model via an expectation-maximization algorithm. After the prediction, the SciSkill is updated by adjusting the SciSkill values based on the difference with the actual game result. Compared to the Elo algorithm, the SciSkill model extracts more detailed information from the matches and describes the player quality more elaborately.

### v3

# football analytics; football scouting; explainable machine learning; player quality; player value; practitioner-oriented research; player development prediction, uncertainty quantification

@@ Abstract
Transfers in professional football (soccer) are risky investments because of the large transfer fees and high risks involved. Although data-driven models can be used to improve transfer decisions, existing models focus on describing players’ historical progress, leaving their future performance unknown. Moreover, recent developments have called for the use of explainable models combined with methods for uncertainty quantification of predictions to improve applicability for practitioners. This paper assesses explainable machine learning models in a practitioner-oriented way for the prediction of the future development in quality and transfer value of professional football players. To this end, the methods for uncertainty quantification are studied through the literature. The predictive accuracy is studied by training the models to predict the quality and value of players one year ahead, equivalent to one season. This is carried out by training them on two data sets containing data-driven indicators describing the player quality and player value in historical settings. In this paper, the random forest model is found to be the most suitable model because it provides accurate predictions as well as an uncertainty quantification method that naturally arises from the bagging procedure of the random forest model. Additionally, this research shows that the development of player performance contains nonlinear patterns and interactions between variables, and that time series information can provide useful information for the modeling of player performance metrics. The resulting models can help football clubs make more informed, data-driven transfer decisions by forecasting player quality and transfer value.
@@
football analytics; football scouting; explainable machine learning; player quality; player value; practitioner-oriented research; player development prediction, uncertainty quantification
[…]
@@ 1 Introduction
Transfers in professional football (soccer) are a risky business because the average transfer fee has increased in recent years (Yang et al., 2024) and because these fees can be characterized as investments with high risks where large fees are involved (McHale and Holmes, 2022). Extensive knowledge about players is beneficial in making well-informed decisions about these complex transfer investments in football. By offering a practitioner-oriented study on forecasting players’ future quality and monetary value, this paper offers methods to football clubs for gaining new insights and for improving their strategic transfer investments.
@@
Models providing information about player quality and value have recently emerged with the evolution of data-driven player performance indicators. Improvements in data-capturing technologies resulted in large data sets containing in-game data about football players, which provide the opportunity to obtain more complex variables on player performance (Rein and Memmert, 2016; Herold et al., 2019). Numerous player performance indicators have been introduced since then. An example is the expected goals (xG) indicator, which values shot chances and shooting ability (Green, 2012; Eggels et al., 2016; Anzer and Bauer, 2021; Mead et al., 2023). Next to such action-specific models, assessment methods exist for general player performance, which can be divided into bottom-up and top-down ratings (Hvattum and Gelade, 2021). Expected threat (xThreat) (Rudd, 2011; Van Roy et al., 2020; Van Arem and Bruinsma, 2024) and VAEP (Decroos et al., 2019; Decroos and Davis, 2020; Van Haaren, 2021; Mendes-Neves et al., 2022) are examples of bottom-up ratings that quantify action quality and use the quality of the actions to create general ratings. Top-down ratings such as plus–minus ratings (Sæbø and Hvattum, 2015; Kharrat et al., 2020; Pantuso and Hvattum, 2021; Hvattum, 2020; De Bacco et al., 2024), Elo ratings adjusted for team sports (Wolf et al., 2020), and the SciSkill algorithm (SciSports, 2020) distribute credit of player performance based on the result of a team as a whole. For the monetary value of players, many models about the estimation of transfer fees and market values have been introduced and provide indicators of the current value of football players (Franceschi et al., 2024). These performance and financial models describe the quality and monetary value of a football player, and they can complement traditional scouting reports. This allows managers and technical directors of football clubs to make better-informed transfer decisions.
@@
These models for player quality and transfer value give information about the quality and financial value of football players up to that moment, although a transfer decision regards whether a football player should be part of a team in the future. To make better-informed transfer decisions, team managers and technical directors also need insights into the future development of the indicator values that describe the financial value and the player performance. This paper examines the training of supervised learning models that forecast the development in player quality and transfer value of players.
@@
To this end, two prediction problems are studied: forecasting the quality of a football player and their transfer value one year ahead. In the first prediction problem, models are trained to predict the development of a top-down quality indicator, the SciSkill (SciSports, 2020). The second prediction problem concerns the prediction of the development of the player value, described by the Estimated Transfer Value (ETV) (SciSports, 2024). A forecasting horizon of one year is selected in this study as it aligns with the length of one full season in football and the success of a transfer is often measured by the performances in the subsequent season. The resulting models of these prediction problems offer insight into the question of whether a player will be worth the money in the future. These models thus provide critical insights for the managerial staff of a professional football club.
@@
To further improve the usability of the research results for staff at football organizations, the findings of models should be presented such that they can be utilized by football clubs in practice, as stressed by Herold et al. (2019). This means that models should not only be assessed on predictive accuracy but also on explainability and on methods for uncertainty quantification (Davis et al., 2024). To this end, only explainable supervised learning models are used in this research, and the models are assessed on their methods for uncertainty quantification. Although deep learning models are common in forecasting tasks, their lack of explainability makes them less suitable for application at football clubs, and they are thus excluded in this study. To focus even more on the applicability for practitioners, the accuracy of the models is partly determined by estimating the loss values on different groups of football players, because certain types of football players are more important from a practitioner’s perspective.
@@
The aim of this research is to find the most suitable explainable machine learning model to forecast player performance with respect to predictive accuracy and methods for uncertainty quantification. To find this, several explainable supervised learning models are studied. By reviewing the literature, favorable models are identified based on their methods for uncertainty quantification. The predictive performance is assessed by implementing them to forecast the player quality (SciSkill) and the player value (Estimated Transfer Value) of football players one year ahead. The predictive quality is studied on both the general population of players and on subgroups of players that are more interesting for practitioners, such as young or high-value players. An overview of this process is visualized in Figure 1. These results will then be combined to find an answer to the research question.
@@
The main contribution of this paper is the illustration of how extra value can be added to football player KPIs by forecasting future values in a practitioner-oriented way. The results of this study provide knowledge on what models are most suitable for application by practitioners, lowering the threshold for real-life implementation. By taking explainability, uncertainty quantification, and performance on important subpopulations of players into account, this research is presented such that the results can be used easily by practitioners. In this way, this paper also contributes to bridging the gap between academic research and practitioners.
This paper is organized as follows: The scientific background in the existing literature is reviewed in Section 2. After a discussion of the considered supervised models and their methods for uncertainty quantification, favorable models are identified based on their methods for uncertainty quantification. Then, the literature on data-driven player value and on quality quantification methods is covered, along with research on forecasting these values. The long-term forecasting of player development is then studied for the two prediction problems, for which the methods are described in Section 3. The results of the models in the prediction problems of the player quality and player value are presented in Section 4. The conclusions of this research are then summarized in Section 5, followed by a discussion of the implications and future directions in Section 6.
@@ 2.1.1 Considered Models
The first linear model considered is ordinary least squares (OLS), sometimes called (multiple) linear regression, as described by (Section 3.2, Hastie et al. (2009)). This model assumes that the true relation is linear $y_{i}=X_{i}\beta+\varepsilon_{i}$, where $\varepsilon_{i}$ is normally distributed noise. OLS minimizes the residual sum of squares
@@
and has the lowest mean square error of all unbiased estimators. Lasso regression Tibshirani (1996) is another linear model that introduces a bias in the model by minimizing the penalized residual sum of squares $RRS(\beta)+\lambda\sum_{i=1}^{p}|\beta_{i}|$. By introducing this bias, it reduces the variance, and it is possible to provide more accurate predictions. The third linear model is the linear mixed effects (LME) model, which assumes that
@@
where $\varepsilon_{i}\sim N(0,\sigma^{2}\Lambda_{i})$ Lindstrom and Bates (1988). The random variable $Z_{i}$ is called the random effect of group $i$ and can be used to model the influence of a specific random attribute that should not influence the predictions, e.g., nationality. Because of their linearity, these models are explainable and suitable for applications.
The CART decision tree (Section 9.2, Hastie et al. (2009)) is a model that recursively splits the feature space by taking splits of the form $\{X|X_{j}\leq s\},\{X|X_{j}>s\}$ such that it minimizes the sum of squares within each split. As described by Hastie et al. (2009) (p. 312), decision trees generally suffer from a high variance. A random forest model solves this by repeatedly training decision trees on resampled data (Chapter 15, Hastie et al. (2009)). The random forest model then predicts by taking the average prediction of these decision trees. Boosting is another way of improving decision trees. A boosting algorithm fits decision trees sequentially and reweighs the data points for which the model has a bad performance (Chapter 10, Hastie et al. (2009)). The XGBoost Chen and Guestrin (2016) combines several techniques, such as regularization and feature quantile estimates, to provide such a boosting algorithm. Generally, the splitting behavior in tree-based algorithms mimics the if–then reasoning of humans. Additionally, they can provide feature importances by considering how much a specific feature improves the predictions on the training set. This makes the tree-based models explainable and thus suitable models.
The last type of model considered in this paper is the k-nearest neighbor (kNN) model. Given the features, these models search for the most similar data points in the training data set and predict the associated $y_{i}$-value by taking the average of the $k$-nearest neighbors (Subsection 2.3.2, Hastie et al. (2009)). The closest neighbors can be weighted more heavily as illustrated by Dudani (1976). The ReliefF algorithm by Robnik-Sikonja and Kononenko (1997) calculates feature importances by looking at the probability of having different values considering the neighbors of a data point. Additionally, the kNN models can provide examples of training data points to explain predictions. This gives an intuitive interpretation to practitioners, since they can see that a player’s forecast performance is based on previous players in similar situations, making the rationale behind the prediction clear and relatable. To interpret a prediction, practitioners can then investigate the players in similar situations, which can help them assess the individual prediction. However, similarity becomes an abstract concept for practitioners in higher dimensions (pp. 108–109, Molnar (2019)), and the kNN model suffers from the curse of dimensionality (Section 2.5, Hastie et al. (2009)), which makes application with a large feature set infeasible. To mitigate this in the current paper, kNN models are applied to a feature set of limited dimensions.
@@ 2.1.2 Uncertainty Quantification
Linear regression has an underlying theory that gives prediction intervals as described by Neter et al. (2004). These prediction intervals are based on assumptions that are unlikely to hold for the prediction problems in this paper, such as the normality of errors. Nonetheless, they do give an indication of the uncertainty of the prediction. Similarly, the kNN models provide the neighbors, which are a group of similar data points. Uncertainty quantification can be obtained by taking the minimal and maximal values of the dependent variable within these neighbors if the number of neighbors $k$ is of significant size. Lastly, the bagging procedure of the random forest model can be utilized to obtain uncertainty quantification for the predictions, as described by Wager et al. (2014). This method utilizes the different decision trees in the random forest to quantify the uncertainty in the prediction. From these properties, it is concluded that the linear regression, random forest, and kNN-based models are favorable with respect to uncertainty quantification.
@@ 2.2 Existing Indicators
In recent years, the increasing amount of available data in football has driven the introduction of methods to rate individual football players (Arntzen and Hvattum, 2021). The performance of a professional football player has traditionally been determined via expert judgment based on video data and statistics describing the frequency of in-game events. Data-driven models have offered the possibility to reduce the bias in assessments and improve the consistency of the judgment of both player quality and player value, as discussed below. In this way, these models have provided new and consistent insights into the quality and monetary value of football players, which has aided in transfer decisions.
@@ 2.2.1 Player Performance
The creation and the comparison of models for player performance have revealed new challenges. Because teams can have different aims in football and can apply various tactics, there does not exist a ground truth for player performance (Davis et al., 2024). A player can, for instance, be instructed to keep the ball in possession, which leads to the player performing fewer actions that might result in scoring a goal. Because of this lack of ground truth, and various models exist that describe different aspects of the game.
Bottom-Up Ratings
Some models define the players’ quality by their actions, called bottom-up ratings Hvattum and Gelade (2021). Although action-specific models exist (Green, 2012; Eggels et al., 2016; Goes et al., 2018; Anzer and Bauer, 2021; Mead et al., 2023), methods to assess the quality of all types of actions have been introduced that generalize the action-specific models. These general models often define a ‘good’ action as one that increases the probability of scoring and decreases the probability of conceding a goal.
@@
The VAEP model by Decroos et al. (2019) calculates the probability of scoring given the last three actions, including in-game context. Because the involved machine learning techniques are considered a black-box model by practitioners, the authors in Decroos and Davis (2020); Van Haaren (2021) introduced methods to make the VAEP model more accessible for practitioners. A comparable, more interpretable framework is the Expected Threat (xThreat) model by Rudd (2011), which only considers the current situation, defined by the location of the ball-possessing player, to estimate the probabilities of the ball transitioning to somewhere else on the pitch. This is modeled by a Markov chain, which can be used to describe the probabilities of scoring before and after each action to find the quality of an action. Van Roy et al. (2020) have compared the xThreat and VAEP models and have found that, although the xThreat model is more interpretable to practitioners, it can only take into account the position of an action and excludes contextual information such as the position of defenders from its model. This means that there is a trade-off between explainability and the inclusion of in-game context when choosing either VAEP or xThreat models. Van Arem and Bruinsma (2024) have extended the xThreat model by including variables describing the defensive situation and height of the ball. This Extended xThreat model can take into account in-game context like the VAEP model, while maintaining explainability as an xThreat model.
These bottom-up ratings describe the quality of a professional football player by assessing in-game on-the-ball actions. These on-the-ball actions are often offensive actions, and they are better at capturing the quality of attackers and attacking midfielders, which leads to a bias when these models are studied. This bias makes it harder to assess the quality of defensive players. Moreover, these models need more detailed event data describing in-game events, limiting the scale on which they can be applied. Consequently, the quality of football players in this study is not assessed with bottom-up ratings.
Top-Down Ratings
@@
These problems of bottom-up ratings Hvattum and Gelade (2021) can be reduced by using top-down models that describe player quality using the lineups and outcomes. Plus–minus ratings are an example of such ratings and were first used in ice hockey and basketball (Kharrat et al., 2020), and were later applied to football by Sæbø and Hvattum (2015). For plus–minus ratings, the game is partitioned into game segments that contain the same lineups, which correspond to the data points in the data set. In this data set, the result of the game segment, the goal difference, for example, is the dependent variable. Indicators describing whether a player was active in the segment are the independent variables. Linear regression is then applied to estimate the influence of players on the results. The coefficients of the regression describe the average impact of a player on the game result and give an indicator for player performance over the period of time covered by the data.
Because substitutions are infrequent in football, a game does not have many game segments with different lineups. Moreover, football is a low-scoring sport. This creates a situation where a low number of segments with limited distinction in outcomes must be used to infer player quality. To deal with this, other quantities have been used as the dependent variable to describe the result of a segment, like the expected number of goals (xG), the expected number of points (xP), and the created VAEP values by Kharrat et al. (2020) and Hvattum and Gelade (2021). Pantuso and Hvattum (2021) additionally showed the potential of taking age, cards, and home advantage into account, and Hvattum (2020) illustrated how separate defensive and offensive ratings can be obtained. Because putting a player in the lineup is an action that a coach performs, De Bacco et al. (2024) adapted the plus–minus ratings using a causal model to better describe the influence of the selection of a player. These studies show how plus–minus ratings have been adapted to the application of rating players in football.
@@
Whereas the plus–minus ratings describe the quality of a player using multiple historical games, there also exist models that determine the quality of a player after each game using the lineups and final score. Elo ratings are such ratings and were originally developed to evaluate performance in one-on-one sports. The concept was subsequently adapted to the game of football by Wolf et al. (2020). This adapted algorithm provides ratings for each individual football player and calculates the team rating via the average of players in a game weighted by the number of minutes played. The ratings are then used to predict the match outcome using a fixed logistic function, and after each game, the individual ratings are adjusted. If the outcome is better than predicted, the player rating is increased, and if the outcome is worse than expected, it is decreased. The authors also introduced an indicator for player impact to deal with the fact that this rating undervalues good players at below-average teams.
The SciSkill (SciSports, 2020) is an industry-validated, more elaborate version of the Elo rating. Instead of only considering one player’s quality, it describes football players combining a defensive rating and an offensive rating. These offensive and defensive scores are then combined to obtain one value, the SciSkill. For each game, the outcome is predicted using a model via an expectation-maximization algorithm. After the prediction, the SciSkill is updated by adjusting the SciSkill values based on the difference with the actual game result as with other Elo-ratings. Compared to the Elo algorithm, which was implemented and validated by Wolf et al. (2020), the SciSkill model extracts more detailed information from the matches and describes the player quality more elaborately.